In an experiment on the photoelectric effect using radiation of wavelength 4.00 x 10^-7m, the maximum electron energy was observed to be 1.40 x 10-19J. With radiation of wavelength 3.00 x 10^-7m the maximum energy was 3.06 x 10^-19J. Derive a value for Planck’s constant.
1
Expert's answer
2020-10-20T07:16:01-0400

Let "K.E_{max}" represent the maximum Kinetic energy of the electron

Let "h" represent Planck's constant

Let "c" represent the speed of light

Let "W_0" represent the work fuction of the metal

Let "\lambda" represent the wavelength of the radiation


"h\frac{c}{
\lambda} - W_0 = K.E_{max}\\
Therefore,\\
h × \frac{3 × 10^8ms^{-1}}{4.00 × 10^{-7}m} - W_0 = 1.40 × 10^{-19}J\\
Also,\\
h × \frac{3 × 10^8ms^{-1}}{3 × 10^{-7}m} - W_0 = 3.06 × 10^{-19}J\\
7.5 × 10^{14}s^{-1}h - W_0 = 1.40 × 10^{-19}J\hspace{1cm}(1)\\
1 × 10^{15}s^{-1}h = 3.06 × 10^{-19}J\hspace{2cm}(2)\\ 
Subtracting\hspace{0.1cm}(1) from\hspace{0.1cm}(2), \\
2.5 × 10^{14}s^{-1}h = 1.66 × 10^{-19}J\\
h = 6.64 × 10^{-34}Js"